HETAV PATEL

Grade 11,

Find: 1% of 1 hour Find: 1% of 1 hour Find: 1% of 1 hour

Find: 
 1% of 1 hour
Find: 
 1% of 1 hour
Find: 
 1% of 1 hour

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We know that, 1 hour = 60 minutes
Then,
1% of 60 minutes
1 minute = 60 seconds
60 minutes = 60 × 60
= 3600 seconds
Now,
1% of 3600 seconds
= (1/100) × 3600
= 1 × 36

Thanks